# Deep Reinforcement Learning for Infinite Horizon Mean Field Problems in Continuous Spaces

### v1

@@ 4.1 Representation of $\mu$ via Score-matching
The question of how to represent and update $\mu$ in the continuous-space setting deserves special consideration in this work. In [Angiuli et al., 2023c, Angiuli et al., 2022b], the authors deal with the discrete-space mean field distribution in a natural way, using a normalized vector containing the probabilities of each state. Each individual state is modeled as a one-hot vector (a Dirac delta measure), and the approximation $\mu_{n}$ is updated at each step using an exponentially weighted update of the form $\mu_{n+1}=\mu_{n}+\rho_{\mu}(\delta_{X_{t_{n}}}-\mu_{n})$ with the mean field learning rate $\rho_{\mu}>0$. [Frikha et al., 2023] uses a similar update in the context of an AC algorithm for solving only MFC problems, while focusing on a more in-depth treatment of the continuous-time aspect. The authors in [Perrin et al., 2021] tackle continuous state spaces for the MFG problem using the method of normalizing flows, which pushes forward a fixed latent distribution, such as a Gaussian, using a series of parameterized invertible maps [Rezende and Mohamed, 2015]. There is reason to believe that other deep generative models, such as generative adversarial networks (GANs) or variational auto-encoders (VAEs), may yield successful representations of the population distribution with their own drawbacks and advantages.
In our case, partly due to its simplicity of implementation, we opt for the method known as score-matching [Hyvärinen, 2005], which has been successfully applied to generative modeling [Song and Ermon, 2019]. If $\mu$ has a density function $p_{\mu}:\mathbb{R}^{d}\to\mathbb{R}$, then its score function is defined as
$s_{\mu}(x)=\nabla\log p_{\mu}(x).$
@@
$L_{\Sigma}(\varphi_{n})\coloneqq\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}.$ (4.5)
@@
$\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}(\varphi_{n})$ (4.6)
@@ 4.2 Unifying Mean Field Game and Mean Field Control Problems
Having laid out the general algorithm, we now address the issue of unifying the MFG and MFC formulations in the style of [Angiuli et al., 2022b, Angiuli et al., 2023c].
@@
The intuitions presented in Section 2 regarding the difference between MFG and MFC suggest that the interplay between the learning rates $\rho_{\Pi}$, $\rho_{V}$, and $\rho_{\Sigma}$ may be used to differentiate between the two solutions of the mean field problem. Taking $\rho_{\Sigma}<\min\{\rho_{\Pi},\rho_{V}\}$ emulates the notion of solving the classical control problem corresponding to a fixed (frozen) $\mu$—or, in this case, a slowly moving $\mu$—and then updating the distribution to match the law of the sate process in an iterative manner. This matches the strategy discussed in Section 2.1 for finding an MFG equilibrium. Conversely, taking $\rho_{\Sigma}>\max\{\rho_{\Pi},\rho_{V}\}$ is more in keeping with simultaneous optimization of the mean field and the policy, which should yield the MFC solution as discussed in Section 2.2.
For a more rigorous justification of the correspondence between the learning rates and mean field problem solutions in the vein of Borkar’s two timescale approach [Borkar, 1997, Borkar, 2008], consult [Angiuli et al., 2022b, Angiuli et al., 2023c, Angiuli et al., 2023d].
@@ 5.1 A Linear-Quadratic Benchmark
We test our algorithm on a 1-dimensional linear-quadratic (LQ) mean field problem where we wish to optimize
@@
Both the MFG and MFC problems corresponding to eqs. 5.1 and 5.2 have explicit analytic solutions, which we state now using the notation consistent with the full derivations in [Angiuli et al., 2022b].
@@ 5.2 Solution for Asymptotic Mean Field Game
Define the constants
$\hat{\Gamma}_{2}=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}\right)}}{4}\qquad\text{and}\qquad\hat{\Gamma}_{1}=-\frac{2\hat{\Gamma}_{2}c_{3}c_{4}}{\hat{\Gamma}_{2}(\beta+2\hat{\Gamma}_{2})-c_{1}c_{2}}.$
@@
$\hat{\alpha}(x)=-\left(2\hat{\Gamma}_{2}x+\hat{\Gamma}_{1}\right).$ (5.3)
@@ 5.3 Solution for Asymptotic Mean Field Control
Proceeding as above, we define the constants
$\Gamma^{*}_{2}=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}\right)}}{4}\qquad\text{and}\qquad\Gamma^{*}_{1}=-\frac{2\Gamma_{2}^{*}c_{3}c_{4}}{\Gamma^{*}_{2}(\beta+2\Gamma^{*}_{2})+c_{5}-c_{1}c_{2}(2-c_{2})}.$
Then the optimal control for the MFC is
$\alpha^{*}(x)=-\left(2\Gamma_{2}^{*}x+\Gamma^{*}_{1}\right).$ (5.6)
@@ 5.4 Hyperparameters and Numerical Specifics
For our numerical experiment, we test our algorithm on two different sets of values for the running cost coefficients $c_{1}$ to $c_{5}$ and volatility $\sigma$ as listed in Tables 3 and 4. The discount factor is fixed in both cases to $\beta=1$, and the continuous time is discretized using step size $\Delta t=0.01$. The critic and score functions are both feedforward neural networks with one hidden layer of 128 neurons and a tanh activation function. The actor is also a feedforward neural network that outputs the mean and standard deviation of a normal distribution from which an action is sampled. Its architecture consists of a shared hidden layer of size 64 neurons and a tanh activation followed by two separate layers of size 64 neurons for the mean and standard deviation. The standard deviation layer is bookended by a softmax activation function to ensure its output is positive. The actor is meant to converge to a deterministic policy—also known as a pure control—over time, so in order to ensure a minimal level of exploration, we add a baseline value of $10^{-5}$ to the output layer. This straightforwardly mimics the notion of entropy regularization detailed in [Wang et al., 2020]. Refer to Table 1 for the learning rates used by the actor, critic, and score networks. Table 2 summarizes the total parameter count for each neural network.
@@
For the Langevin Monte Carlo iterations, we pick a step size $\epsilon=5\times 10^{-2}$ as shown in Table 1. Rather than beginning the iterations at the $n^{th}$ step with samples $x_{0}=(x_{0}^{(1)},x_{0}^{(2)},\dots,x_{0}^{(k)})$ from an arbitrary distribution, we take $x_{0}=S_{t_{n-1}}$, the samples generated from the Langevin dynamics in the previous step, to accelerate convergence. We run 200 iterations at each step using $k=1000$ samples.
The results of the algorithm applied to the LQ benchmark problem after $N=10^{6}$ iterations are displayed in figs. 2 and 4 with different sets of parameters along with the corresponding analytic solutions. We observe many of the same insights alluded to by [Angiuli et al., 2023c, Angiuli et al., 2022b] regarding the differences in recovering the MFG versus the MFC solution. Specifically, convergence to the MFG solution is more stable and faster than convergence to the MFC solution, as evidenced by the convergence plots in figs. 2 and 4. Further, in both cases, there were certain runs in which instability was amplified by the AC algorithm, in which case we saw the weights of the neural networks diverge to numerical overflow. In order to combat this, we imposed a bound on the state space during the first 200,000 iterations, truncating all states to the interval $[-5,5]$. We removed the artificial truncation following the initial iterations and were able to mitigate the instability issues leading to overflow.
@@
Table 1: Choice of learning rates used to obtain results in all figures seen in this work—$\rho_{\Pi}$ for the actor, $\rho_{V}$ for the critic, $\rho_{\Sigma}$ for the score function, and $\epsilon$ for the Langevin dynamics. Boldface values indicate a difference in the learning rate between the MFG and MFC regimes.
MFG MFC
$\rho_{\Pi}$ $5\times 10^{-6}$ $5\times 10^{-6}$
$\rho_{V}$ $10^{-5}$ $10^{-5}$
$\rho_{\Sigma}$ $\mathbf{10^{-6}}$ $\mathbf{5\times 10^{-4}}$
$\epsilon$ $5\times 10^{-2}$ $5\times 10^{-2}$
@@
Figure 1: The histogram (grey) is the learned asymptotic distribution using samples generated from the parameterized score function $\Sigma_{\varphi_{N}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
Table 3: Running cost coefficients and volatility for eqs. 5.1 and 5.2. The results for this parameter set are displayed in figs. 2 and 2. $c_{1}$ $c_{2}$ $c_{3}$ $c_{4}$ $c_{5}$ $\sigma$ 0.25 1.5 0.5 0.6 1.0 0.3 Refer to caption (a) Result for $\rho_{\Sigma}=10^{-6}$ Refer to caption (b) Result for $\rho_{\Sigma}=5\times 10^{-4}$
Refer to caption (a) Results for $\rho_{\Sigma}=10^{-6}$ Refer to caption (b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
Figure 1: The histogram (grey) is the learned asymptotic distribution using samples generated from the parameterized score function $\Sigma_{\varphi_{N}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
Figure 2: The blue curve is a rolling average of the absolute error between the mean of samples produced from the parameterized score function $\Sigma_{\varphi_{n}}$ and the optimal mean $\hat{m}$ from eq. 5.5 in the case of MFG (left) and $m^{*}$ from eq. 5.8 in the case of MFC (right). Large jumps are due to random outliers which result from the stochasticity of our algorithm.
Figure 3: The histogram (grey) is the learned asymptotic distribution using samples generated from $\Sigma_{\varphi_{n}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.
@@
Table 4: Running cost coefficients and volatility for eqs. 5.1 and 5.2. The results for this parameter set are displayed in figs. 4 and 4. $c_{1}$ $c_{2}$ $c_{3}$ $c_{4}$ $c_{5}$ $\sigma$ 0.15 1.0 0.25 1.0 2.0 0.5 Refer to caption (a) MFG results for $\rho_{\Sigma}=10^{-6}$ Refer to caption (b) MFC results for $\rho_{\Sigma}=5\times 10^{-4}$
Refer to caption (a) Results for $\rho_{\Sigma}=10^{-6}$ Refer to caption (b) Results for $\rho_{\Sigma}=5\times 10^{-4}$
@@
Figure 3: The histogram (grey) is the learned asymptotic distribution using samples generated from $\Sigma_{\varphi_{n}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The $x$-axis shows the state variable $x$, the left $y$-axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu(x)$.

### v3

@@ 4.1 Representation of $\mu$ via Score-matching
The question of how to represent and update $\mu$ in the continuous-space setting deserves special consideration in this work. In [Angiuli et al., 2023b, Angiuli et al., 2022b], the authors deal with the discrete-space mean field distribution in a natural way, using a normalized vector containing the probabilities of each state. Each individual state is modeled as a one-hot vector (a Dirac delta measure), and the approximation $\mu_{n}$ is updated at each step using an exponentially weighted update of the form $\mu_{n+1}=\mu_{n}+\rho_{\mu}(\delta_{X_{t_{n}}}-\mu_{n})$ with the mean field learning rate $\rho_{\mu}>0$. [Frikha et al., 2023] uses a similar update in the context of an AC algorithm for solving only MFC problems, while focusing on a more in-depth treatment of the continuous-time aspect. The authors in [Perrin et al., 2021] tackle continuous-state spaces for the MFG problem using the method of normalizing flows, which pushes forward a fixed latent distribution, such as a Gaussian, using a series of parameterized invertible maps [Rezende and Mohamed, 2015]. There is reason to believe that other deep generative models, such as generative adversarial networks (GANs) or variational auto-encoders (VAEs), may yield successful representations of the population distribution with their own drawbacks and advantages.
In our case, partly due to its simplicity of implementation, we opt for the method known as score-matching [Hyvärinen, 2005], which has been successfully applied to generative modeling [Song and Ermon, 2019]. If $\mu$ has a density function $p_{\mu}:\mathbb{R}^{d}\to\mathbb{R}$, then its Stein score function is defined as
$s_{\mu}(x)\coloneqq\nabla\log p_{\mu}(x).$
@@
$L_{\Sigma}^{(n)}(\varphi_{n})\coloneqq\operatorname{tr}\left(\nabla_{x}\Sigma_{\varphi_{n}}(X_{t_{n}})\right)+\frac{1}{2}\left\lVert\Sigma_{\varphi_{n}}(X_{t_{n}})\right\rVert_{2}^{2}.$ (4.5)
@@
$\varphi_{n+1}=\varphi_{n}-\rho_{\Sigma}\nabla_{\varphi}L_{\Sigma}^{(n)}(\varphi_{n})$ (4.6)
@@ 4.2 Unifying Mean Field Game and Mean Field Control Problems
Having laid out the general algorithm, we now address the issue of unifying the MFG and MFC formulations in the style of [Angiuli et al., 2022b, Angiuli et al., 2023b].
@@
The intuitions presented in Section 2 regarding the difference between MFG and MFC suggest that the interplay between the learning rates $\rho_{\Pi}$, $\rho_{V}$, and $\rho_{\Sigma}$ may be used to differentiate between the two solutions of the mean field problem. Taking $\rho_{\Sigma}<\min\{\rho_{\Pi},\rho_{V}\}$ emulates the notion of solving the classical control problem corresponding to a fixed (frozen) $\mu$—or, in this case, a slowly moving $\mu$—and then updating the distribution to match the law of the state process in an iterative manner. This matches the strategy discussed in Section 2.1 for finding an MFG equilibrium. Conversely, taking $\rho_{\Sigma}>\max\{\rho_{\Pi},\rho_{V}\}$ is more in keeping with simultaneous optimization of the mean field and the policy, which should yield the MFC solution as discussed in Section 2.2.
Borkar’s stochastic multi-scale approximation framework [Borkar, 1997, Borkar, 2008] has been applied in [Angiuli et al., 2022b, Angiuli et al., 2023b, Angiuli et al., 2023c] to describe how the choice of learning rates is crucial to derive the solution of different mean field problems using a 2-time scale Q-learning algorithm. Following similar ideas, we show how our method can be expressed as a 3-time scale actor-critic algorithm converging to MFG or MFC solution depending on the choice of learning rates.
In particular, Algorithm 1 is characterized by the following system of numerical updates
$\begin{cases}\quad\varphi_{n+1}&=\varphi_{n}-\rho_{n}^{\Sigma}\nabla_{\varphi}L_{\Sigma}(\varphi_{n}),\\ \quad\theta_{n+1}&=\theta_{n}-\rho_{n}^{V}\nabla_{\theta}L_{V}(\theta_{n}),\\ \quad\psi_{n+1}&=\psi_{n}-\rho_{n}^{\Pi}\nabla_{\psi}L_{\Pi}(\psi_{n}),\end{cases}$ (4.7)
where $L_{\Sigma}$, $L_{V}$, and $L_{\Pi}$ (each of which actually depend implicitly on all three of the parameter vectors) are the loss functions of the Stein score, the critic, and the actor networks, respectively, with $\rho_{n}^{\Sigma}$, $\rho_{n}^{V}$ and $\rho_{n}^{\Pi}$ being the corresponding learning rates satisfying usual Robbins–Monro type conditions, namely: $\sum_{n=0}^{\infty}\rho_{n}^{\Sigma}=\sum_{n=0}^{\infty}\rho_{n}^{V}=\sum_{n=0}^{\infty}\rho_{n}^{\Pi}=\infty$ and $\sum_{n=0}^{\infty}{(\rho_{n}^{\Sigma})}^{2}=\sum_{n=0}^{\infty}{(\rho_{n}^{V})}^{2}=\sum_{n=0}^{\infty}{(\rho_{n}^{\Pi})}^{2}<\infty$
The relationship between learning rates is crucial to determine the convergence of our algorithm. As shown by [Borkar and Konda, 1997], the actor-critic paradigm can be represented as a multi-scale stochastic approximation procedure in which the updates of the actor evolve at a much slower pace than the ones of the critic. Indeed, as discussed in Section 3.2.1, choosing $\rho_{\Pi}<\rho_{V}$ allows the actor $\Pi$ to be seen as frozen by the critic $V$ which can be learned accordingly. Our algorithm extends the traditional 2-time scale framework of actor-critic algorithms to approach mean field problems by adding an additional time scale for learning the mean field distributions.
Three time scale approach for MFG.
If $\rho^{\Sigma}_{n}<\rho^{\Pi}_{n}<\rho^{V}_{n}$ so that $\rho^{\Sigma}_{n}/\rho^{\Pi}_{n}\mapsto 0$ and $\rho^{\Pi}_{n}/\rho^{V}_{n}\mapsto 0$ as $n\mapsto\infty$, the system 4.7 tracks the ODE system:
$\begin{cases}\quad\dot{\varphi}&=-\nabla_{\varphi}L_{\Sigma}(\varphi,\psi,\theta)\\ \quad\dot{\psi}&=-\frac{1}{\epsilon}\nabla_{\psi}L_{\Pi}(\varphi,\psi,\theta)\\ \quad\dot{\theta}&=-\frac{1}{\epsilon\tilde{\epsilon}}\nabla_{\theta}L_{V}(\varphi,\psi,\theta)\end{cases}$ (4.8)
where $\rho^{\Sigma}_{n}/\rho^{\Pi}_{n}$ and $\rho^{\Pi}_{n}/\rho^{V}$ are thought of being of order $\epsilon\ll 1$ and $\tilde{\epsilon}\ll 1$, respectively. In the case $\varphi$ and $\psi$ are fixed, we assume the ODE $\dot{\theta}=-\frac{1}{\epsilon\tilde{\epsilon}}\nabla_{\theta}L_{V}(\varphi,\psi,\theta)$ to have a stable equilibrium $(\varphi,\psi,\theta^{\varphi,\psi})$ corresponding to the local minimum of the operator $L_{V}$, meaning that the network approximates the value function $V$ corresponding to the policy $\Pi$ and the population distribution $\mu$ described respectively by the parameters $\psi$ and $\varphi$. Similarly, given fixed $\varphi$, we assume the ODE $\dot{\psi}=-\frac{1}{\epsilon}\nabla_{\psi}L_{\Pi}(\varphi,\psi,\theta_{\varphi,\psi})$ to reach a stable equilibrium $(\varphi,\psi_{\varphi},\theta_{\varphi,\psi_{\varphi}})$ as the local minimum of the operator $L_{\Pi}$, meaning that the network approximates the optimal policy of the infinitesimal player facing the crowd distribution $\mu$ with Stein score function parameterized by $\varphi$. Finally, the ODE $\dot{\varphi}=-\nabla_{\varphi}L_{\Pi}(\varphi,\psi_{\varphi},\theta_{\varphi,\psi_{\varphi}})$ stabilizes at $(\varphi^{*},\psi_{\varphi^{*}},\theta_{\varphi^{*},\psi_{\varphi^{*}}})$, the local minimum of $L_{\Sigma}$ which results the network approximating the Stein score function of the population at equilibrium of the MFG problem.
Three time scale approach for MFC.
If $\rho^{\Pi}_{n}<\rho^{V}_{n}<\rho^{\Sigma}_{n}$ so that $\rho^{\Pi}_{n}/\rho^{V}_{n}\mapsto 0$ and $\rho^{V}_{n}/\rho^{\Sigma}_{n}\mapsto 0$ as $n\mapsto\infty$, the system 4.7 tracks the ODE system:
$\begin{cases}\quad\dot{\psi}&=-\nabla_{\psi}L_{\Pi}(\varphi,\psi,\theta)\\ \quad\dot{\theta}&=-\frac{1}{\epsilon}\nabla_{\theta}L_{V}(\varphi,\psi,\theta)\\ \quad\dot{\varphi}&=-\frac{1}{\epsilon\tilde{\epsilon}}\nabla_{\varphi}L_{\Sigma}(\varphi,\psi,\theta)\end{cases}$ (4.9)
where $\rho^{\Pi}_{n}/\rho^{V}_{n}$ and $\rho^{V}_{n}/\rho^{\Sigma}$ are thought of being of order $\epsilon\ll 1$ and $\tilde{\epsilon}\ll 1$ respectively. Considering $\psi$ and $\theta$ to be fixed, we assume the ODE $\dot{\varphi}=-\frac{1}{\epsilon\tilde{\epsilon}}\nabla_{\varphi}L_{\Sigma}(\varphi,\psi,\theta)$ to have a stable equilibrium $(\varphi_{\psi,\theta},\psi,\theta)$ corresponding to the local minimum of the operator $L_{\Sigma}$, meaning that the network approximates the population distribution induced by the policy and value function described respectively by the parameter $\psi$ and $\theta$. Similarly, fixed $\psi$, we assume the ODE $\dot{\theta}=-\frac{1}{\epsilon}\nabla_{\theta}L_{V}(\varphi_{\psi,\theta},\psi,\theta)$ to reach a stable equilibrium $(\varphi_{\psi,\theta_{\psi}},\psi,\theta_{\psi})$ as the local minimum of the operator $L_{V}$, meaning that the network approximates the value function corresponding to the policy $\Pi$ described by $\psi$. Finally, the ODE $\dot{\psi}=-\nabla_{\psi}L_{\Pi}(\varphi_{\psi,\theta_{\psi}},\psi,\theta_{\psi}))$ stabilizes at $(\varphi_{\psi^{*},\theta_{\psi^{*}}},\psi^{*},\theta_{\psi^{*}})$, the local minimum of $L_{\Pi}$ which corresponds to the network approximating the optimal control strategy of the MFC problem.
Writing a rigorous proof of convergence with precise assumptions on these operators is extremely complicated and outside of the scope of this paper. It was recently achieved in the case of Q-learning with finite-state and -action spaces in [Angiuli et al., 2023c].
@@ 5.1 A Linear-Quadratic Benchmark
We test our algorithm on a linear-quadratic (LQ) mean field problem where we wish to optimize
@@
Both of the 1-dimensional MFG and MFC problems corresponding to eqs. 5.1 and 5.2 have explicit analytic solutions, which we state now using the notation consistent with the derivations in [Angiuli et al., 2022b]. We also test our algorithm on the equivalent multivariate problem for which we include the full derivation of the solutions in Appendix A.
@@ 5.2 Solution for Asymptotic Mean Field Game
Traditional methods for deriving the LQ problem solution begin with recovering the value function
$v(x)\coloneqq\inf_{\alpha\in\mathbb{A}}\mathbb{E}\left[\left.\int_{0}^{\infty}e^{-\beta t}\left(\frac{1}{2}\alpha_{t}^{2}+c_{1}\left(X_{t}-c_{2}m\right)^{2}+c_{3}\left(X_{t}-c_{4}\right)^{2}+c_{5}{m}^{2}\right)\,\mathrm{d}t\right\rvert\,X_{0}=x\right]$
as the solution of a Hamilton-Jacobi-Bellman equation. In the MFG case, we denote the optimal value function as $\hat{v}$ and an explicit form is given by $\hat{v}(x)=\hat{\Gamma}_{2}x^{2}+\hat{\Gamma}_{1}x+\hat{\Gamma}_{0}$, where
$\hat{\Gamma}_{2}=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}\right)}}{4}$
$\hat{\Gamma}_{1}=-\frac{2\hat{\Gamma}_{2}c_{3}c_{4}}{\hat{\Gamma}_{2}(\beta+2\hat{\Gamma}_{2})-c_{1}c_{2}}$
$\hat{\Gamma}_{0}=\frac{c_{5}\hat{m}^{2}+c_{3}c_{4}^{2}+c_{1}c_{2}^{2}\hat{m}^{2}+\sigma^{2}\hat{\Gamma}_{2}-\frac{1}{2}\hat{\Gamma}_{1}^{2}}{\beta}.$
@@
$\hat{\alpha}(x)=-\hat{v}^{\prime}(x)=-\left(2\hat{\Gamma}_{2}x+\hat{\Gamma}_{1}\right).$ (5.3)
@@ 5.3 Solution for Asymptotic Mean Field Control
Similarly as above, we denote the MFC value function by $v^{*}$ and claim that is has the form $v^{*}(x)=\Gamma^{*}_{2}x^{2}+\Gamma_{1}^{*}x+\Gamma^{*}_{0}$ with
$\Gamma^{*}_{2}=\frac{-\beta+\sqrt{\beta^{2}+8\left(c_{1}+c_{3}\right)}}{4}$
$\Gamma^{*}_{1}=-\frac{2\Gamma_{2}^{*}c_{3}c_{4}}{\Gamma^{*}_{2}(\beta+2\Gamma^{*}_{2})+c_{5}-c_{1}c_{2}(2-c_{2})}$
$\Gamma^{*}_{0}=\frac{c_{5}{m^{*}}^{2}+c_{3}c_{4}^{2}+c_{1}c_{2}^{2}{m^{*}}^{2}+\sigma^{2}\Gamma^{*}_{2}-\frac{1}{2}{\Gamma^{*}_{1}}^{2}}{\beta}.$
The optimal control for the MFC is
$\alpha^{*}(x)=-{v^{*}}^{\prime}(x)=-\left(2\Gamma_{2}^{*}x+\Gamma^{*}_{1}\right).$ (5.6)
@@
5.4 A multivariate Linear-Quadratic Benchmark
We also test our algorithm on a higher-dimensional generalization of the scalar LQ problem in eqs. 5.2 and 5.1 with the cost functional
$\mathbb{E}\left[\int_{0}^{\infty}e^{\beta t}\left(\frac{1}{2}\alpha_{t}^{T}\alpha_{t}+(x-C_{2}m)^{T}C_{1}(x-C_{2}m)+(x-c_{4})^{T}C_{3}(x-c_{4})+m^{T}C_{5}m\right)\,\mathrm{d}t\right],$ (5.9)
and state dynamics
$\mathrm{d}X_{t}=\alpha_{t}\,\mathrm{d}t+\sigma\,\mathrm{d}W_{t},\qquad t\in[0,\infty).$ (5.10)
In this case, $C_{1},C_{3},C_{5}\in\mathbb{R}^{d\times d}$ are positive-definite matrices, $c_{4}\in\mathbb{R}^{d}$, $C_{2}\in\mathbb{R}^{d\times d}$, $\sigma\in\mathbb{R}^{d\times m}$, and $W_{t}$ is an $m$-dimensional Brownian motion. We test our algorithm on a benchmark problem for the case $d=m=2$, and refer the reader to Appendix A for the derivations of the analytic solutions to both the MFG and MFC problems associated with the above system.
@@ 5.5 Hyperparameters and Numerical Specifics
For our numerical experiment, we test our algorithm on two different sets of values for the running cost coefficients and volatility $\sigma$ as listed in Tables 3 and 4 for the 1-dimensional case and Table 7 for the 2-dimensional case. The discount factor is fixed in all cases to $\beta=1$, and the continuous time is discretized using step size $\Delta t=0.01$. The critic and score functions are both feedforward neural networks with one hidden layer of 128 neurons and a tanh activation function. For the 1-dimensional examples, the actor is also a feedforward neural network that outputs the mean and standard deviation of a normal distribution from which an action is sampled. Its architecture consists of a shared hidden layer of size 64 neurons and a tanh activation followed by two separate layers of size 64 neurons for the mean and standard deviation. The standard deviation layer is bookended by a softmax activation function to ensure its output is positive. The actor is meant to converge to a deterministic policy—also known as a pure control—over time, so in order to ensure a minimal level of exploration, we add a baseline value of $10^{-5}$ to the output layer. This straightforwardly mimics the notion of entropy regularization detailed in [Wang et al., 2020]. The actor for the 2-dimensional example uses the same number of hidden layers and hidden neurons, and, correspondingly, outputs a mean vector and a diagonal covariance matrix to induce exploration. Refer to Table 1 for the learning rates used by the actor, critic, and score networks which were selected via grid search to give the best results. Table 2 summarizes the total parameter count for each neural network.
@@
For the Langevin Monte Carlo iterations, we pick a step size $\epsilon=5\times 10^{-2}$ as shown in Table 1, and we run 200 iterations at each step using $k=1000$ samples.
The results of the algorithm applied to the LQ benchmark problem after $N=10^{6}$ iterations are displayed in figs. 2, 4, 7, 9, 12, 13 and 16 with different sets of parameters along with the corresponding analytic solutions. In addition to the plots of the learned solutions, we also provide plots for various error values to observe convergence with respect to the number of training steps. We consider the following error metrics for the MFG problem and their corresponding counterparts for the MFC problem:
  • The absolute error in the learned population mean:
    $e_{\hat{m}}(n)\coloneqq||m_{t_{n}}-\hat{m}||_{2},$
  • The expected absolute error in the learned control:
    $e_{\hat{\alpha}}(n)\coloneqq\mathbb{E}_{x\sim\hat{\mu}}\left[||\alpha_{\psi_{n}}(x)-\hat{\alpha}(x)||_{2}\right]\approx\frac{1}{\ell}\sum_{j=1}^{\ell}||\alpha_{\psi_{n}}(x_{j})-\hat{\alpha}(x_{j})||_{2},$
  • The expected absolute error in the learned value function:
    $e_{\hat{v}}(n)\coloneqq\mathbb{E}_{x\sim\hat{\mu}}\left[||V_{\theta_{n}}(x)-\hat{v}(x)||_{2}\right]\approx\frac{1}{\ell}\sum_{j=1}^{\ell}||V_{\theta_{n}}(x_{j})-\hat{v}(x_{j})||_{2},$
where $x_{j}$ are i.i.d. with distribution $\hat{\mu}$ for $j=1,2,\dots,\ell$.
@@
Table 1: Choice of learning rates used to obtain results in all figures seen in this work—$\rho_{\Pi}$ for the actor, $\rho_{V}$ for the critic, $\rho_{\Sigma}$ for the score function, and $\epsilon$ for the Langevin dynamics. The two left columns are the learning rates used for the case $d=1$ and the right two for $d=2$. Boldface values indicate a difference in the learning rate between the MFG and MFC regimes.
$d=1$ $d=2$
MFG MFC MFG MFC
$\rho_{\Pi}$ $5\times 10^{-6}$ $5\times 10^{-6}$ $5\times 10^{-6}$ $5\times 10^{-6}$
$\rho_{V}$ $10^{-5}$ $10^{-5}$ $10^{-5}$ $10^{-5}$
$\rho_{\Sigma}$ $\mathbf{10^{-6}}$ $\mathbf{5\times 10^{-4}}$ $\mathbf{10^{-6}}$ $\mathbf{10^{-5}}$
$\epsilon$ $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$ $5\times 10^{-2}$
@@
It should be noted that taking larger values of the volatility $\sigma$ yields degrading results in the learning of the solution as a result of the lower signal-to-noise ratio in the state samples. Tables 6 and 6 illustrate the increase in several error metrics at the final time step $N$ for the 1-dimensional experiment with all other values fixed.
We observe many of the same insights alluded to by [Angiuli et al., 2023b, Angiuli et al., 2022b] regarding the differences in recovering the MFG versus the MFC solution. Specifically, convergence to the MFG solution is more stable and faster than convergence to the MFC solution, as evidenced by the convergence plots in figs. 2, 7, 12, 5, 10, 15, 4, 9 and 15. Further, in all cases, there were certain runs in which instability was amplified by the AC algorithm, in which case we saw the weights of the neural networks diverge to numerical overflow. In order to combat this, we imposed a bound on the state space during the first 200,000 iterations, truncating all states to the interval $[-5,5]$ for 1-dimension case and the box $[-5,5]\times[-5,5]$ in the 2-dimensional case. We removed the artificial truncation following the initial iterations and were able to mitigate the instability issues leading to overflow.
@@
Table 3: Running cost coefficients and volatility for eqs. 5.1 and 5.2. The results for this parameter set are displayed in figs. 2, 2, 4, 4 and 5.
$c_{1}$ $c_{2}$ $c_{3}$ $c_{4}$ $c_{5}$ $\sigma$
0.25 1.5 0.5 0.6 1.0 0.3
(a) Result for $\rho_{\Sigma}=10^{-6}$
(b) Result for $\rho_{\Sigma}=5\times 10^{-4}$
@@
Figure 1: The histogram (blue) is the learned asymptotic distribution using samples generated from the parameterized score function $\Sigma_{\varphi_{N}}$ and the dashed line (blue) is the learned feedback control after $N=10^{6}$ iterations averaged over five runs with different initial samples. The green curves correspond to the optimal control and mean field distribution for MFC, while the orange curves are the equivalent for MFG. The bottom axis shows the state variable $x$, the left axis refers to the value of the control $\alpha(x)$, and the right axis represents the probability density of $\mu$.
(c) Results for $\rho_{\Sigma}=10^{-6}$
(d) Results for $\rho_{\Sigma}=5\times 10^{-4}$